### v1

# Can GPT-O1 Kill All Bugs?

Abstract
ChatGPT has long been proven to be effective in automatic program repair (APR). With the continuous iterations and upgrades of the ChatGPT version, its performance in terms of fixes has already reached state-of-the-art levels. However, there are few works comparing the effectiveness and variations of different versions of ChatGPT on APR. In this work, we evaluate the performance of the latest version of ChatGPT (O1-preview and O1-mini), ChatGPT-4o, and historical version of ChatGPT on APR. We study the improvements of the O1 model over traditional ChatGPT in terms of APR from multiple perspectives (repair success rate, repair cost, behavior patterns), and find that O1’s repair capability exceeds that of traditional ChatGPT, successfully fixing all 40 bugs in the benchmark. Our work can serve as a reference for further in-depth exploration of the applications of ChatGPT in APR.
Can GPT-O1 Kill All Bugs?
Haichuan Hu1, Ye Shang2, Guolin Xu3, Congqing He4, Quanjun Zhang2 thanks:  Corresponding author. 1Alibaba Cloud 2State Key Laboratory for Novel Software Technology, Nanjing University, China 3School of Computer Science and Technology, Chongqing University, China 4University Sains Malaysia huhaichuan.hhc@alibaba-inc.com; {quanjun.zhang,ye.shang}@smail.nju.edu.cn; xuguolin0626@gmail.com; hecongqing@hotmail.com;
1 Introduction
With the development of automated program repair (APR) technology, researchers have proposed many effective methods to fix bugs automatically. Traditional template-based APR Xia et al. (2023); Xia and Zhang (2022) performs well for known bugs but are powerless against previously unseen bugs. Learning-based APR Chen et al. (2022); Chi et al. (2022) is able to learn the bug-fixing patterns automatically from existing code repositories, thus having good generalization capability. As a further development of learning-based APR, large language model (LLM) based APR methods are receiving increasing attention.
ChatGPT OpenAI (2023), which is widely regarded as the most representative application of LLMs, has proven to perform well across various code-related tasks Sun et al. (2023); Liu et al. (2023); Yuan et al. (2023). Moreover, ChatGPT has been proven to be effective in program repair as well. Sobania et al.  Sobania et al. (2023) evaluate the bug-fixing capabilities of ChatGPT and find ChatGPT is able to fix 31 out of 40 bugs on QuixBugs Lin et al. (2017) benchmark. Xia et al. Xia and Zhang (2023) employ ChatGPT in a conversational manner to fix 114 and 48 bugs on the Defects4J-v1.2 and Defects4J-v2.0 benchmark.
However, few studies have compared the effectiveness and differences of various versions of ChatGPT in APR tasks. Most people understand ChatGPT in terms of the fact that newer versions have more parameters and larger training datasets. As a result, it can better understand code, leading to improved performance in APR. Such perspective may be challenged with occurrence of the latest version of ChatGPT, ChatGPT-o1 (O1). Different from previous versions of ChatGPT, O1 leverages reinforcement learning (RL) and chain of thought (COT) techniques. Before answering a question, it first takes time to think, organizes a chain of thought, and then arrives at a final answer. By practice, O1 is more suitable for fields with complex logic and relatively definitive answers, such as mathematics and programming.
In our work, we evaluates O1’s performance in APR through dialogues, given the limited understanding of O1. We conduct experiments on the two currently released trial versions of the O1 model (O1-mini and O1-preview), and GPT-4o as a representative of the most popular version of ChatGPT. We design a two-step repair process. First, we provide a basic prompt template for ChatGPT to directly repair the bugs. For bugs that fail the test cases, we further provide ChatGPT with the error information from the test cases to perform a second round of repair. After all repairs are done, we collect data on repair success rate, time spent on repairs, and behavior patterns of the model during the repair process, to further analyze O1’s performance in APR. The results show that O1 outperforms baseline (31/40) and ChatGPT-4o (38/40) in APR, successfully repairing all 40 bugs in the Quixbugs benchmark. Moreover, O1’s unique chain of thought reasoning pattern has proven effective for APR tasks. By forming a chain of thought, O1 can better understand the logic of the code, provide repair ideas, and deliver correct repair code.
2 Methodology
In this section, we present the method for evaluating ChatGPT’s ability to fix bugs. We use different versions of ChatGPT to fix benchmark bugs and evaluate the repair behavior from multiple dimensions.
2.1 Two-step Fix
The process of bug fixing by ChatGPT consists of two steps. First, we use a basic repair template to ask the model whether there are bugs in the target code snippet. After making a judgment, the model is asked to fix the bugs.
Listing 1: A basic template used to fix program bugs.
Does this program have a bug? How to fix
it?
[Code of the Buggy Program]
Listing 1 shows a basic template we use to fix program bugs. We begin the prompt with the question Does this program have a bug? How to fix it? to request ChatGPT to fix the following python program, leaving a blank line between the question and the program.
After ChatGPT is asked to fix the buggy program, it will provide an initial fix answer. Previous versions of ChatGPT (such as ChatGPT-4o and earlier) would directly provide an answer, while the newer versions of ChatGPT (ChatGPT-o1) first spend some time thinking, develop a chain of thought, and then fix the program based on the chain of thought before presenting the complete fix code.
We verify the correctness of the answer provided by ChatGPT with existing test set. For buggy programs that do not pass tests, we perform a secondary fix. To help ChatGPT better understand the reasons for the program errors, we include the error messages in the template used for the secondary fix. Listing 2 is the prompt template used in the secondary fix step. Following The given corrected version fails to pass the test cases, and the results are as follows:, we provide the original error report to prompt ChatGPT for further repair.
Listing 2: The secondary fix template for bug fixing.
The given corrected version fails to
pass the test cases, and the results are
as follows:
==== FAILURES ====
[Error Message of Test Cases]
2.2 Evaluate O1 and Other ChatGPT models
We evaluate ChatGPT’s performance in fixing program bugs from the following dimensions.
  • Repair success rate, which is the most important metric for evaluating the effectiveness of ChatGPT in program repair.
  • Time for thinking. Compared to earlier versions of ChatGPT, O1 has added a "thinking" phase. We record and analyze the time spent by O1-preview and O1-mini in thinking phase.
  • Response length. We record and compare the output lengths of different models to evaluate the monetary cost of repairing.
  • Model behavior pattern. We analyze behavior patterns of O1 and other ChatGPT models to explore their mind when fixing bugs.
3 Experiment and Results
3.1 Benchmark and Baselines
We use QuixBugs Lin et al. (2017) as our benchmark. For each of the 40 benchmark problems, we take the erroneous python version. These programs have complete context and corresponding test cases, and they are relatively short, making them suitable for bug fixing through dialogue with ChatGPT. We compare our results with previous work Sobania et al. (2023). In prompt designing, we keep consistent with previous work Sobania et al. (2023) and carefully review the comments in the original program to ensure that they do not reveal the solution, retaining only the relevant parts of the test cases. Thus, we integrate the baseline method into the first step of the repair method presented in Section 2.1.
3.2 Results
In this section, we first compare the repair results of the latest ChatGPT model O1, previous versions of ChatGPT, and the baseline method on QuixBugs. We then analyze and summarize the details of the O1 model’s repair behavior.
3.3 Comparison of O1, Previous Versions of ChatGPT, and Baseline Results
Table 1: Results achieved by O1-preview, O1-mini, GPT-4o, ChatGPT, Codex Prenner et al. (2022), CoCoNut Lutellier et al. (2020), and the standard APR approaches Ye et al. (2021) on the problems from the QuixBugs benchmark. For ChatGPT, the number of successful runs are listed in brackets. The baseline method Sobania et al. (2023) also mentions that when additional information is given, the repair effectiveness of ChatGPT improves. We highlight these results in blue. For O1-preview, O1-mini and GPT-4o, we provide the failed test case output in the prompt and ask the model to try fixing it again. We mark the results of the second fix in red.
Benchmark problem    O1-preview    O1-mini    GPT-4o ChatGPT     Codex     CoCoNut     Standard APR
bitcount ✗(0 / 4)
breadth-first-search ✓(2 / 4)
bucketsort ✓(4 / 4)
depth-first-search ✗(0 / 4)
detect-cycle ✗(0 / 4)
find-first-in-sorted ✓(2 / 4)
find-in-sorted ✓(3 / 4)
flatten ✓(4 / 4)
gcd ✗(0 / 4)
get-factors ✓(1 / 4)
hanoi ✗(0 / 4)
is-valid-parenthesization ✓(2 / 4)
kheapsort ✗(0 / 4)
knapsack ✓(1 / 4)
kth ✗(0 / 4)
lcs-length ✗(0 / 4)
levenshtein ✗(0 / 4)
lis ✗(0 / 4)
longest-common-subsequence ✗(0 / 4)
max-sublist-sum ✗(0 / 4)
mergesort ✓(1 / 4)
minimum-spanning-tree ✗(0 / 4)
next-palindrome ✓(1 / 4)
next-permutation ✗(0 / 4)
pascal ✓(1 / 4)
possible-change ✓(1 / 4)
powerset ✗(0 / 4)
quicksort ✓(1 / 4)
reverse-linked-list ✓(2 / 4)
rpn-eval ✗(0 / 4)
shortest-path-length ✓(1 / 4)
shortest-path-lengths ✗(0 / 4)
shortest-paths ✓(1 / 4)
shunting-yard ✓(2 / 4)
sieve ✗(0 / 4)
sqrt ✓(1 / 4)
subsequences ✓(1 / 4)
to-base ✗(0 / 4)
topological-ordering ✗(0 / 4)
wrap ✗(0 / 4)
$\Sigma$ (Solved) 38(40) 37(40) 35(38) 19(31) 21 19 7
We present the main results of comparison in Table 1. In terms of baseline results, despite the significant improvement in repair effectiveness after introducing additional information, with a repair success rate reaching 31/40, it is still far from the new version of ChatGPT (ChatGPT-4o, O1-mini, O1-preview). It can be seen that ChatGPT has made significant improvements in program repair capabilities through iterations.
Compared to the current mainstream version of ChatGPT (ChatGPT-4o), O1 shows some improvement in program repair capabilities. Before providing test case error information, O1-mini and O1-preview can repair 37 and 38 bugs respectively, which is 2 and 3 more than ChatGPT-4o. After providing test case error information, both O1-mini and O1-preview are able to repair all 40 bugs, whereas ChatGPT-4o can only repair 38 bugs.
To further investigate the improvements of the O1 model compared to previous versions of ChatGPT in program repair, we also conduct a case analysis on the programs that the O1 model is able to repair but ChatGPT-4o and previous versions of ChatGPT can not. We find that these programs are relatively complex, often involving structures like recursion and nested loops, and they are likely to correspond to solutions for real-world problems rather than simple utility methods (such as bitcount). Take hanoi as instance, both O1-preview and O1-mini successfully repair the bug on the first attempt, whereas neither ChatGPT-4o nor the baseline version of ChatGPT are able to fix it. In the baseline method, ChatGPT fails four times before finally succeeding after being provided with the correct test case answers. With only error information and no test case answers, ChatGPT-4o is still unable to repair it correctly. ChatGPT-4o incorrectly interpret the boundary condition, assuming that the source rod cannot be equal to the destination rod. In order to solve the hanoi problem, the O1 model spends 15 seconds thinking and forms a chain of thought (analyze functionality, check code logic, correct steps, optimize move steps). This chain of thought helps it correctly understand the logic of the problem, avoiding falling into incorrect logical branches. From this, we can infer that the reasoning pattern from chain of thought to solution is crucial for repairing buggy programs with more complex logic.
3.4 Evaluation of O1 in Terms of Response Time, Response Length, and Behavior Pattern
In terms of response length, we conduct a token count analysis of O1-preview, O1-mini, and ChatGPT-4o on all benchmark programs using the tool tiktoken. We found that the average response lengths for the three models are 1450, 1086, and 654 tokens, respectively. This indicates that the O1 model’s response length is over 50% longer compared to previous versions of ChatGPT, leading to more comprehensive and complete answers, albeit at a higher cost.
In terms of response time, since the generation time is related to the length of the output, we only measure the time taken by the O1 model for thinking. We found that the average thinking time for O1-preview reaches 19.82 seconds, which is about three times longer than that of O1-mini (7.02 seconds).
In terms of model behavior patterns, we observe that the O1 model typically begins by performing logical analysis on the buggy program, generating a solution, and then gradually proceeding with repairs and testing before providing the complete fixed code. In contrast, ChatGPT-4o tends to provide the repaired code first, followed by an explanation of the code. Please refer to Appendix A for detailed evaluatin.
4 Conclusion
This paper evaluates the capabilities of the latest ChatGPT model O1 in program repair. By comparing it with previous versions of ChatGPT and baseline results, the paper demonstrates the advancements of the O1 model. Additionally, we conduct case analysis and conclude the advantages of the O1 model in repairing more complex bugs. The paper also provides a comprehensive analysis of other characteristics of the O1 model, including thinking time, response length, and behavior patterns, offering a reference for future in-depth research on utilizing O1 for program repair.
5 Limitations
Currently, O1 is in the trial phase, with usage limits and high API costs. Understanding of O1 is still quite insufficient. The benchmark used in our study is relatively small, and further research is needed with larger datasets to explore the O1 model’s capabilities in APR more comprehensively.
Appendix A Detailed Evaluation of O1
Table 2: The token length output by O1 and ChatGPT-4o when fixing each benchmark problem.
Benchmark Problem O1-preview O1-mini ChatGPT-4o
bitcount 1846 1172 647
breadth-first-search 1179 1015 780
bucketsort 1540 710 669
depth-first-search 1620 738 667
detect-cycle 1518 719 757
find-first-in-sorted 2099 1406 820
find-in-sorted 2159 912 785
flatten 987 856 487
gcd 1393 822 456
get-factors 1109 932 553
hanoi 1826 1262 726
is-valid-parenthesization 878 578 770
kheapsort 1282 1203 1078
knapsack 1722 1029 904
kth 1848 1171 472
lcs-length 1345 927 589
levenshtein 1485 914 727
lis 1612 896 597
longest-common-subsequence 1299 875 710
max-sublist-sum 1423 1370 485
mergesort 575 1017 1438
minimum-spanning-tree 1386 1346 903
next-palindrome 1215 1057 1658
next-permutation 1642 931 817
pascal 1778 798 692
possible-change 940 860 798
powerset 1642 1128 563
quicksort 760 691 281
reverse-linked-list 1062 1005 342
rpn-eval 1638 1097 415
shortest-path-length 814 1097 768
shortest-path-lengths 2035 1263 432
shortest-paths 1130 1244 495
shunting-yard 2104 1329 573
sieve 1432 1067 511
sqrt 931 812 268
subsequences 1557 653 363
to-base 968 1015 281
topological-ordering 1497 2755 390
wrap 1374 752 383
Average 1450.07 1086.90 654.07
Table 3: The thinking time spent on repairing each benchmark problem using the O1 model. When thinking time is less than five seconds, we count it as five seconds.
Benchmark problem O1-preview O1-mini ChatGPT-4o
bitcount 17s < 5s ~
breadth-first-search 27s < 5s ~
bucketsort 18s < 5s ~
depth-first-search 9s < 5s ~
detect-cycle 23s < 5s ~
find-first-in-sorted 31s 11s ~
find-in-sorted 18s < 5s ~
flatten 19s < 5s ~
gcd 8s < 5s ~
get-factors 16s < 5s ~
hanoi 15s 7s ~
is-valid-parenthesization 10s < 5s ~
kheapsort 30s 6s ~
knapsack 38s 5s ~
kth 28s 10s ~
lcs-length 12s < 5s ~
levenshtein 18s < 5s ~
lis 51s 38s ~
longest-common-subsequence 15s 42s ~
max-sublist-sum 21s 6s ~
mergesort 17s < 5s ~
minimum-spanning-tree 28s 10s ~
next-palindrome 53s 8s ~
next-permutation 23s 7s ~
pascal 33s 8s ~
possible-change 14s < 5s ~
powerset 23s < 5s ~
quicksort 11s < 5s ~
reverse-linked-list 14s < 5s ~
rpn-eval 16s 5s ~
shortest-path-length 16s 7s ~
shortest-path-lengths 13s < 5s ~
shortest-paths 21s < 5s ~
shunting-yard 17s < 5s ~
sieve 15s 8s ~
sqrt 10s 5s ~
subsequences 47s < 5s ~
to-base 20s < 5s ~
topological-ordering 18s 6s ~
wrap 21s < 5s ~
Average 19.82s 7.02s ~